OpenStudy (anonymous):

would the answer be 64 ft? H (t ) = -16t + 64t 14.Find the maximum height of the ball. (Hint: this is the maximum function value)

OpenStudy (campbell_st):

well is the equation \[H(t) = -16t^2 + 64t\]

OpenStudy (anonymous):

probably + 14 you need to complete the square to have your parabola in the form: \[y = a(x-h)^2 + k\] the point (x, y) = (h, k) is the vertex. if a < 0 [our case] (h, k) is the highest point if a > 0 (h, k) is the lowest point

OpenStudy (campbell_st):

if so... the max height is on the line of symmetry of the equation \[t = -\frac{b}{2a}\] you have a = -16 and b = 64 find t and substitute to find max height

OpenStudy (jdoe0001):

the maximum height will be found at the vertex of the parabola equation, and for an equation like so \(\bf ax^2+bx+c\) you can get the vertex coordinates at \(\bf \left(-\cfrac{b}{2a}\quad ,\quad c-\cfrac{b^2}{4a}\right)\)

OpenStudy (campbell_st):

or you can find the 1st derivative and the solve that... for t. then substitute to solution into the original equation...

OpenStudy (ranga):

The answer you got is correct.

OpenStudy (anonymous):

oh go0d! thank you